mathslover (mathslover):

if : \(\large{\frac{1}{x+1}+\frac{2}{y+2}+\frac{2006}{z+2006}=1}\) find the value of : \(\large{\frac{x^2}{x^2+x}+\frac{y^2}{y^2+2y}+\frac{z^2}{z^2+2006z}}\)

mathslover (mathslover):

@ash2326 , @ajprincess @mathavraj

mathslover (mathslover):

I did like this : \[\large{(\frac{x}{x+1})+(\frac{y}{y+2})+(\frac{z}{z+2006})}\]

mathslover (mathslover):

but what to do next?:

OpenStudy (dominusscholae):

Do long division for each fraction.

OpenStudy (dominusscholae):

\[x/(x+1) = 1-(1/x+1)\]

OpenStudy (dominusscholae):

\[2/(y+2) = 1-2/(y+2)\]

OpenStudy (ganpat):

 does it make sense ??